# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ 5.1 WDM Particle Mass Posteriors
To contextualize these results, note that Nadler et al. (2020b) infer $M_{\mathrm{hm}}<10^{7.3}~{}M_{\mathrm{\odot}}$ ($m_{\mathrm{WDM}}>7.9~{}\mathrm{keV}$) at $68\%$ confidence from current DES and PS1 observations of the MW satellite population and simultaneously disfavor the “strong” cutoff scenario. Our forecasts are consistent with these bounds and indicate that observations of the remaining MW satellite population alone can improve WDM particle mass constraints by $\approx 60\%$ in the “weak” cutoff scenario. These projected one-host WDM constraints do not significantly change if we use MW-like hosts that contain Large Magellanic Cloud (LMC) analog subhalos; however, modeling the actual MW satellite population without accounting for LMC-associated satellites can bias inferred WDM constraints (e.g., see the discussion in Nadler et al. 2021b).
@@
In Scenario C, $M_{\mathrm{hm}}$ is only probed at the $\approx 1\sigma$ level using one host, mainly due to a significant degeneracy with $\mathcal{M}_{50}$. When two complete hosts are used, the non-detection of $M_{\mathrm{hm}}$ persists while the $1\sigma$ uncertainty on $\mathcal{M}_{50}$ shrinks. In our model, $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ both suppress satellite abundances as a function of peak halo mass, and their observable effects only differ in the shape of the halo mass-dependent suppression imprinted on satellite populations. The non-detection of $M_{\mathrm{hm}}$ is consistent with our finding in Section 4.2 that the shape of the galaxy formation cutoff is difficult to measure, even when multiple hosts are combined. Note that we vary the shape of the galaxy formation cutoff but not of the WDM cutoff, which allows $\mathcal{M}_{50}$ to be disentangled from $\mathcal{S}_{\mathrm{gal}}$ in the inference; it will be interesting to explore constraints on the functional dependence of SHMF suppression further in future work.
@@ 5.2 Projections for Additional Hosts
Following Section 4.3, we perform idealized tests with all galaxy–halo connection parameters held fixed to derived projections for WDM constraints with $N_{\mathrm{hosts}}>2$. Specifically, we calculate idealized likelihoods as a function of $M_{\mathrm{hm}}$ that sample over galaxy–halo connection model realizations and host combinations in the “strong” and “weak” cutoff scenarios, assuming CDM subhalo abundances. We use these to derive projected $2\sigma$ upper limits on $M_{\mathrm{hm}}$, and we convert these to lower limits on $m_{\mathrm{WDM}}$ using Equation 4. Figure 7 shows the results of this test for the “strong” cutoff. We find that statistically-limited measurements of $32$ complete hosts yields $m_{\mathrm{WDM}}\gtrsim 30~{}\mathrm{keV}$ at the $2\sigma$ level; different host combinations contribute a ±$\pm 10~{}\mathrm{keV}$ uncertainty to this projection. These results do not significantly change in the “weak” cutoff scenario because galaxy–halo connection parameters are fixed; in Section 5.1, our $68\%$ confidence WDM limits were more stringent for the “weak” versus the “strong” cutoff due to less prominent galaxy–halo connection degeneracies.
Projections for the scaling of WDM constraints with $N_{\mathrm{hosts}}$ can be derived from the functional dependence of SHMF suppression on $m_{\mathrm{WDM}}$ and Poisson statistics. In particular, given that our mock dwarf galaxy population measurements probe the SHMF down to $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$, we use Equation 5 to predict how lower limits on $m_{\mathrm{WDM}}$, denoted $m_{\mathrm{WDM,lim}}$, depend on $N_{\mathrm{hosts}}$. Assuming that uncertainties on the SHMF amplitude shrink as $N_{\mathrm{hosts}}^{-1/2}$ and that SHMF suppression at $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$ determines the WDM constraint, we derive $m_{\mathrm{WDM,lim}}\propto N_{\mathrm{hosts}}^{3/20}$. Fixing the normalization of this calculation to the result of our idealized test for $N_{\mathrm{hosts}}=4$, we project $2\sigma$ lower limits of
@@
which is valid for $N_{\mathrm{hosts}}\geq 4$ and statistically-limited measurements. Figure 7 confirms this scaling in our idealized tests.
The relatively shallow scaling between $m_{\mathrm{WDM,lim}}$ and $N_{\mathrm{hosts}}$ in Equation 11 follows because, as $m_{\mathrm{WDM}}$ increases, the SHMF rapidly approaches that in CDM over the range of $M_{\mathrm{peak}}$ probed by satellite galaxies. Specifically, the (sub)halos that host dwarf galaxies in our forecasts span a particular range of peak virial masses from $\approx 10^{8}$ to $10^{10}~{}M_{\mathrm{\odot}}$. According to Equation 11, detecting all satellites of the several hundred MW-mass systems within $\approx 40~{}\mathrm{Mpc}$ (e.g., Geha et al. 2017; Mao et al. 2021)—which may be enabled by future optical/infrared space telescopes such as the Habitable Worlds Observatory; for potential capabilities, see the LUVOIR report (The LUVOIR Team 2019)—would constrain models with $m_{\mathrm{WDM}}\approx 40~{}\mathrm{keV}$. Additional dwarf galaxy observables and probes of lower-mass halos below the galaxy formation threshold will be necessary to qualitatively improve this WDM sensitivity.
@@ 5.3 Interpretation
Our results indicate that future dwarf galaxy surveys can potentially place stringent constraints on WDM, comparable to the upcoming sensitivity of other small-scale structure probes (Drlica-Wagner et al., 2019). However, we emphasize that there are significant degeneracies between the effects of galaxy formation and DM physics on dwarf galaxy populations, which must be modeled accurately to for unbiased inference. Dwarf galaxy observables beyond those considered here, including star formation histories (which can be delayed in WDM-like models; e.g., see Bozek et al. 2019) and stellar velocity dispersions (e.g., Kim & Peter 2021; Esteban et al. 2023) may help break these degeneracies. In parallel, combining dwarf galaxy populations with independent probes of low-mass halo abundances, such as strong lensing and stellar streams, is another promising path forward (e.g., see Enzi et al. 2021; Banik et al. 2021a; Nadler et al. 2021b).
Figure 7: Projected $2\sigma$ lower limits on $m_{\mathrm{WDM}}$ in the “strong” cutoff scenario for statistically-limited measurements of dwarf satellite populations, with all galaxy–halo connection parameters held fixed. Points show mean lower limits as a function of the number of hosts combined in the inference for $N_{\mathrm{hosts}}=1$ (blue), $2$ (red), $4$ (orange), $8$ (green), $16$ (cyan), and $32$ (gray); the gray band shows the $68\%$ range of these constraints due to different host combinations. The black dashed line shows the $N_{\mathrm{hosts}}^{3/20}$ scaling expected from the functional dependence of SHMF suppression on $m_{\mathrm{WDM}}$.
Our MCMC forecasts’ sensitivity to WDM masses of $\approx 10$ to $20~{}\mathrm{keV}$ can be understood in terms of the corresponding SHMF suppression. In particular, the left panel of Figure 2 shows $\approx 20\%$ to $50\%$ suppression of the SHMF at peak virial masses of $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$ in the most extreme WDM models that are marginally consistent with our mock CDM data. This is consistent with the fact that our CDM forecasts are sensitive to a “strong” galaxy formation cutoff over the range of peak halo masses indicated by the blue band. Interestingly, our most optimistic constraint of $m_{\mathrm{WDM}}>17.0~{}\mathrm{keV}$ in Scenario B corresponds to SHMF suppression of only $\approx 10\%$, highlighting the importance of breaking additional galaxy–halo connection degeneracies in this case.
@@ 6.1 SHMF Amplitude Posteriors
As summarized in Table 1, we find that SHMF suppression can be constrained to $\approx 70\%$, $60\%$, and $50\%$ that in CDM at peak virial masses of $10^{8}~{}M_{\mathrm{\odot}}$, $10^{9}~{}M_{\mathrm{\odot}}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively, for one complete host in the “weak” cutoff scenario. Meanwhile, enhanced SHMFs are only constrained relative to CDM at the factor of $\approx 20$, $4$, and $3$ levels in these mass decades. Our analysis of the full posteriors in Appendix C.3 reveals degeneracies between SHMF amplitude and galaxy–halo connection parameters including the faint-end luminosity function slope $\alpha$ and luminosity scatter $\sigma_{M}$, particularly at the low-mass end ($M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$). These degeneracies are expected because variations in the underlying SHMF can mimic or counteract the effects of our galaxy–halo connection model for the faintest satellite we consider. At higher masses ($M_{\mathrm{peak}}\approx 10^{10}~{}M_{\mathrm{\odot}}$), we forecast that the SHMF can be measured reasonably precisely.
@@
A complete census of satellites around two hosts improves the precision of these constraints. For example, at $M_{\mathrm{peak}}=10^{10}~{}M_{\mathrm{\odot}}$, $68\%$ confidence limits on SHMF enhancement change from $3$ to $2$ times that in CDM, and limits on SHMF suppression change from $50\%$ to $80\%$ that in CDM; this level of improvement is similar at lower halo masses (see Table 1). Furthermore, with two complete hosts, both the galaxy formation cutoff and underlying SHMF are recovered more confidently at the $1\sigma$ level. Note that our projected SHMF enhancement constraints are not very sensitive to the assumed galaxy formation cutoff; however, SHMF suppression constraints are slightly weaker in our “strong” cutoff scenario because suppression in the underlying SHMF mimics the assumed galaxy formation cutoff. In the one-host case, these degeneracies substantially degrade the measurement of $\mathcal{M}_{50}$ relative to our galaxy formation cutoff forecasts that assume CDM subhalo abundances in Section 4. These results underscore the importance of combining multiple satellite populations, particularly when assumptions about underlying subhalo populations are relaxed.
@@
Figure 8 illustrates the results of our one and two-host SHMF forecasts in the “strong” cutoff scenario relative to the SHMF from our MW-mass zoom-in simulation suite. In the $M_{\mathrm{peak}}=10^{9}~{}M_{\mathrm{\odot}}$ and $10^{10}~{}M_{\mathrm{\odot}}$ decades, our forecasted SHMF constraints are comparable to the host-to-host scatter about the mean SHMF among our forty zoom-in simulations (see Mao et al. 2015 for a detailed discussion of the host-to-host scatter). At $M_{\mathrm{peak}}=10^{8}~{}M_{\mathrm{\odot}}$, SHMF suppression is probed at a level comparable to the host-to-host scatter, while SHMF enhancement that significantly exceeds the range of CDM predictions is compatible with the mock data.
@@
Figure 8: Mean and $16$th–$84$th percentile host-to-host scatter of the SHMF (black line and band) from our forty CDM MW-mass zoom-in simulations. Blue circles (red squares) and error bars show the mean and $68\%$ confidence intervals from our SHMF forecast using one (two) complete satellite populations of MW-mass hosts, assuming a “strong” galaxy formation cutoff. One and two-host error bars are offset horizontally for clarity.
@@
These results imply that secondary properties beyond host halo mass that correlate with SHMF amplitude (e.g., concentration; Mao et al. 2015) can be probed using multiple satellite populations. On the other hand, accurately interpreting SHMF measurements from future dwarf galaxy surveys will require modeling these secondary host halo properties, particularly for models that suppress the SHMF at low masses. We emphasize that our forecasted upper limits on the SHMF are extremely uncertain at the low-mass end; the precision of these upper limits is mainly determined by galaxy–halo connection degeneracies for two or more hosts. Thus, we expect that surveys of dwarfs in a range of environments at the bright end of the satellite populations we consider here will reduce galaxy–halo connection uncertainties and help determine the SHMF.
@@ 6.2 Projections for Additional Hosts
We derive projected uncertainties on $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ by calculating the likelihood of each parameter for fixed galaxy–halo connection parameters in the “strong” cutoff scenario. For large $N_{\mathrm{hosts}}$, these projected SHMF constraints shrink according to $N_{\mathrm{hosts}}^{-1/2}$, as expected from Poisson statistics; furthermore, they do not change significantly in the “weak” cutoff scenario. We project that statistically-limited measurements of $4$ ($32$) complete hosts yield SHMF uncertainties of $\approx 5\%$ ($\approx 2\%$) relative to CDM. Unlike our results in Section 6.1, these projected uncertainties are roughly symmetric and do not depend strongly on peak halo mass because they are not limited by galaxy–halo connection degeneracies.

### v2

@@ 5.1 WDM Particle Mass Posteriors
To contextualize these results, note that Nadler et al. (2020b) inferred $M_{\mathrm{hm}}<10^{7.3}~{}M_{\mathrm{\odot}}$ ($m_{\mathrm{WDM}}>7.9~{}\mathrm{keV}$) at $68\%$ confidence from current DES and PS1 observations of the MW satellite population and simultaneously disfavor the “strong” cutoff scenario. Our forecasts are consistent with these bounds and indicate that observations of the remaining MW satellite population alone can improve WDM particle mass constraints by $\approx 60\%$ in the “weak” cutoff scenario. These projected one-host WDM constraints do not significantly change if we use MW-like hosts that contain Large Magellanic Cloud (LMC) analog subhalos; however, modeling the actual MW satellite population without accounting for LMC-associated satellites can bias inferred WDM constraints (e.g., see the discussion in Nadler et al. 2021b).
@@
In Scenario C, our $M_{\mathrm{hm}}$ posterior is consistent with the input value, indicating that the WDM extension of our inference framework is also self-consistent. We find that $M_{\mathrm{hm}}$ is only probed at the $\approx 1\sigma$ level using one host, mainly due to a significant degeneracy with $\mathcal{M}_{50}$. When two complete hosts are used, the nondetection of $M_{\mathrm{hm}}$ persists while the $1\sigma$ uncertainty on $\mathcal{M}_{50}$ shrinks. In our model, $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ both suppress satellite abundances as a function of peak halo mass, and their observable effects only differ in the shape of the halo mass-dependent suppression imprinted on satellite populations. The nondetection of $M_{\mathrm{hm}}$ is consistent with our finding in Section 4.2 that the shape of the galaxy formation cutoff is difficult to measure, even when multiple hosts are combined. Note that we vary the shape of the galaxy formation cutoff but not of the WDM cutoff, which allows $\mathcal{M}_{50}$ to be disentangled from $\mathcal{S}_{\mathrm{gal}}$ in the inference; it will be interesting to explore constraints on the functional dependence of SHMF suppression further in future work.
@@ 5.2 Projections for Additional Hosts
Following Section 4.3, we perform idealized tests with all galaxy–halo connection parameters held fixed to derived projections for WDM constraints with $N_{\mathrm{hosts}}>2$. Specifically, we calculate idealized likelihoods as a function of $M_{\mathrm{hm}}$ that sample over galaxy–halo connection model realizations and host combinations in the “strong” and “weak” cutoff scenarios, assuming CDM subhalo abundances. We use these to derive projected $2\sigma$ upper limits on $M_{\mathrm{hm}}$, and we convert these to lower limits on $m_{\mathrm{WDM}}$ using Equation 4. Figure 7 shows the results of this test for the “strong” cutoff. We find that statistically limited measurements of $32$ complete hosts yield $m_{\mathrm{WDM}}\gtrsim 30~{}\mathrm{keV}$ at the $2\sigma$ level; different host combinations contribute a ±$\pm 10~{}\mathrm{keV}$ uncertainty to this projection. These results do not significantly change in the “weak” cutoff scenario because galaxy–halo connection parameters are fixed; in Section 5.1, our $68\%$ confidence WDM limits were more stringent for the “weak” versus the “strong” cutoff due to less prominent galaxy–halo connection degeneracies.
We derive projections for the scaling of WDM constraints with $N_{\mathrm{hosts}}$ from the dependence of the SHMF suppression on $m_{\mathrm{WDM}}$. In particular, given that our mock dwarf galaxy population measurements probe the SHMF down to $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$, we use Equation 5 to predict how lower limits on $m_{\mathrm{WDM}}$, denoted $m_{\mathrm{WDM,lim}}$, depend on $N_{\mathrm{hosts}}$. Assuming that uncertainties on the SHMF amplitude shrink as $N_{\mathrm{hosts}}^{-1/2}$ and that SHMF suppression at $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$ determines the WDM constraint, we derive $m_{\mathrm{WDM,lim}}\propto N_{\mathrm{hosts}}^{3/20}$. Fixing the normalization of this calculation to the result of our idealized test for $N_{\mathrm{hosts}}=4$, we project $2\sigma$ lower limits of
@@
which is valid for $N_{\mathrm{hosts}}\geq 4$ and statistically limited measurements. Figure 7 confirms this scaling in our idealized tests.
The relatively shallow scaling between $m_{\mathrm{WDM,lim}}$ and $N_{\mathrm{hosts}}$ in Equation 11 follows because, as $m_{\mathrm{WDM}}$ increases, the SHMF rapidly approaches that in CDM over the range of $M_{\mathrm{peak}}$ probed by satellite galaxies. Specifically, the (sub)halos that host dwarf galaxies in our forecasts span a particular range of peak virial masses from $\approx 10^{8}$ to $10^{10}~{}M_{\mathrm{\odot}}$. According to Equation 11, detecting all satellites of the several hundred MW-mass systems within $\approx 40~{}\mathrm{Mpc}$ (e.g., Geha et al. 2017; Mao et al. 2021)—which may be enabled by future optical/infrared space telescopes such as the Habitable Worlds Observatory (for potential capabilities, see the LUVOIR report; The LUVOIR Team 2019)—would constrain models with $m_{\mathrm{WDM}}\approx 40~{}\mathrm{keV}$. Additional dwarf galaxy observables and probes of lower-mass halos below the galaxy formation threshold will be necessary to qualitatively improve this WDM sensitivity.
@@ 5.3 Interpretation
Our results indicate that future dwarf galaxy surveys can potentially place stringent constraints on WDM, comparable to the upcoming sensitivity of other small-scale structure probes (Drlica-Wagner et al., 2019). However, we emphasize that there are significant degeneracies between the effects of galaxy formation and DM physics on dwarf galaxy populations, which must be modeled accurately to for unbiased inference. Dwarf galaxy observables beyond those considered here, including star formation histories (which can be delayed in WDM-like models; e.g., Bozek et al. 2019) and stellar velocity dispersions (e.g., Kim & Peter 2021; Esteban et al. 2023) may help break these degeneracies. In parallel, combining dwarf galaxy populations with independent probes of low-mass halo abundances, such as strong lensing and stellar streams, is another promising path forward (e.g., Banik et al. 2021a; Nadler et al. 2021b; Enzi et al. 2021).
Figure 7: Projected $2\sigma$ lower limits on $m_{\mathrm{WDM}}$ in the “strong” cutoff scenario for statistically limited measurements of dwarf satellite populations, with all galaxy–halo connection parameters held fixed. Points show mean lower limits as a function of the number of hosts combined in the inference for $N_{\mathrm{hosts}}=1$ (blue), $2$ (red), $4$ (orange), $8$ (green), $16$ (cyan), and $32$ (gray); the gray band shows the $68\%$ range of these constraints due to different host combinations. The black dashed line shows the $N_{\mathrm{hosts}}^{3/20}$ scaling expected from the functional dependence of SHMF suppression on $m_{\mathrm{WDM}}$.
[…]
@@ 6.1 SHMF Amplitude Posteriors
In both the “strong” and “weak” cutoff scenarios, our $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ posteriors are consistent with CDM, validating our inference framework when applied to generalized SHMFs. Furthermore, as summarized in Table 1, we find that SHMF suppression can be constrained to $\approx 70\%$, $60\%$, and $50\%$ that in CDM at peak virial masses of $10^{8}~{}M_{\mathrm{\odot}}$, $10^{9}~{}M_{\mathrm{\odot}}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively, for one complete host in the “weak” cutoff scenario. Meanwhile, enhanced SHMFs are only constrained relative to CDM at factors of $\approx 20$, $4$, and $3$ levels in these mass decades. Uniformly suppressed (enhanced) SHMFs corresponding to the $68\%$ confidence limits from these “weak” cutoff forecasts are shown by the dotted red (dashed blue) lines in the right panel of Figure 2.
@@
Our analysis of the full posteriors in Appendix C.3 reveals degeneracies between SHMF amplitude and galaxy–halo connection parameters including the faint-end luminosity function slope $\alpha$ and luminosity scatter $\sigma_{M}$, particularly at the low-mass end ($M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$). These degeneracies are expected because variations in the underlying SHMF can mimic or counteract the effects of our galaxy–halo connection model for the faintest satellite we consider. At higher masses ($M_{\mathrm{peak}}\approx 10^{10}~{}M_{\mathrm{\odot}}$), we forecast that the SHMF can be measured reasonably precisely.
@@
A complete census of satellites around two hosts improves the precision of these constraints. For example, at $M_{\mathrm{peak}}=10^{10}~{}M_{\mathrm{\odot}}$, $68\%$ confidence limits on SHMF enhancement change from three to two times that in CDM, and limits on SHMF suppression change from $50\%$ to $80\%$ that in CDM; this level of improvement is similar at lower halo masses (see Table 1). Furthermore, with two complete hosts, both the galaxy formation cutoff and underlying SHMF are recovered more confidently at the $1\sigma$ level. Note that our projected SHMF enhancement constraints are not very sensitive to the assumed galaxy formation cutoff; however, SHMF suppression constraints are slightly weaker in our “strong” cutoff scenario because suppression in the underlying SHMF mimics the assumed galaxy formation cutoff. In the one-host case, these degeneracies substantially degrade the measurement of $\mathcal{M}_{50}$ relative to our galaxy formation cutoff forecasts that assume CDM subhalo abundances in Section 4. These results underscore the importance of combining multiple satellite populations, particularly when assumptions about underlying subhalo populations are relaxed.
@@
Figure 8 illustrates the results of our one- and two-host SHMF forecasts in the “strong” cutoff scenario relative to the SHMF from our MW-mass zoom-in simulation suite. In the $M_{\mathrm{peak}}=10^{9}~{}M_{\mathrm{\odot}}$ and $10^{10}~{}M_{\mathrm{\odot}}$ decades, our forecasted SHMF constraints are comparable to the host-to-host scatter about the mean SHMF among our $40$ zoom-in simulations (see Mao et al. 2015 for a detailed discussion of the host-to-host scatter). At $M_{\mathrm{peak}}=10^{8}~{}M_{\mathrm{\odot}}$, SHMF suppression is probed at a level comparable to the host-to-host scatter, while SHMF enhancement that significantly exceeds the range of CDM predictions is compatible with the mock data.
@@
Figure 8: Mean and $16$th–$84$th percentile host-to-host scatter of the SHMF (black line and band) from our $40$ CDM MW-mass zoom-in simulations. Blue circles (red squares) and error bars show the mean and $68\%$ confidence intervals from our SHMF forecast using one (two) complete satellite populations of MW-mass hosts, assuming a “strong” galaxy formation cutoff. For clarity, the one- and two-host error bars are offset horizontally.
@@
These results imply that secondary properties beyond host halo mass that correlate with SHMF amplitude (e.g., concentration; Mao et al. 2015) can be probed using multiple satellite populations. On the other hand, accurately interpreting SHMF measurements from future dwarf galaxy surveys will require modeling these secondary host halo properties, particularly for models that suppress the SHMF at low masses. We emphasize that our forecasted upper limits on the SHMF are very weak at low masses; the precision of these upper limits is mainly determined by galaxy–halo connection degeneracies for two or more hosts. Thus, we expect surveys of dwarfs in a range of environments, at the bright end of the satellite populations we consider here, to reduce galaxy–halo connection uncertainties and help determine the SHMF.
@@ 6.2 Projections for Additional Hosts
We derive projected uncertainties on $\xi_{8}$, $\xi_{9}$, and $\xi_{10}$ by calculating the likelihood of each parameter for fixed galaxy–halo connection parameters in the “strong” cutoff scenario. For large $N_{\mathrm{hosts}}$, these projected SHMF constraints shrink according to $N_{\mathrm{hosts}}^{-1/2}$, as expected from Poisson statistics; furthermore, they do not change significantly in the “weak” cutoff scenario. We project that statistically limited measurements of $4$ ($32$) complete hosts yield SHMF uncertainties of $\approx 5\%$ ($\approx 2\%$) relative to CDM. Unlike our results in Section 6.1, these projected uncertainties are roughly symmetric and do not depend strongly on peak halo mass because they are not limited by galaxy–halo connection degeneracies.